OpenStudy (anonymous):

Evaluate -4(a+b) - 10a/b when a=8 and b=5

terenzreignz (terenzreignz):

\[\Large -4(\color{red}a+\color{blue}b) - \frac{10\color{red}a}{\color{blue}b} \] \[\large \color{red}{a=8\\\\\large \color{blue}{b=5}}\] First step? Replace all a's with 8 and all b's with 5.

OpenStudy (anonymous):

It's 0.675

OpenStudy (anonymous):

\[-4(8+5)-10\frac{ 8 }{ 5 } = -32-20-\frac{ 80 }{ 5 } = -68 Ans. -68\]

OpenStudy (anonymous):

thank you both!!! i feel so dumb

OpenStudy (anonymous):

No worries :)